100. Sailboat speed. The sail area- displaced ratio S provides a measure of the sail power available to drive the boat. For a boat with a displacement of d pounds and a sail area A square feet
S=16Ad^-2/3
a. Find S to the nearest tenth for the Tartan 4100, which has a sail area of 810 square feet and a displacement of 23,245 pounds.
b. Solve the formula for d.

Answer by stanbon(75887)

S=16Ad^-2/3
a. Find S to the nearest tenth for the Tartan 4100, which has a sail area of 810 square feet and a displacement of 23,245 pounds.
S=16(810)(23,245)^(-2/3)
S=12960*0.0012277604
S=15.9117752
------------------------
b. Solve the formula for d.
d^(-2/3)=S/(16A)
d=[S/16A]^(-3/2)
---------------
Cheers,
Stan H.